Consider the triangle OAB where O=(0,0),B=(3,4).If orthocentre of the triangle is H(1,4),the coordinate of A is?
 

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Product of slopes of perpendicular lines =-1

Let A be (x,y)
Slope AH *slope OB = -1

3x+4y= 19

Slope AB*slope OH = -1

x+4y= 19

Solving 2 eqn

x=0
y = 19/4

Regards